OpenStudy (anonymous):

A bag contains 10 white golf balls and 6 striped golf balls. A golfer wants to add 112 golf balls to the bag. He wants the ratio of white to striped golf balls to remain the same. How many of each should he add?

OpenStudy (anonymous):

plz can u help me

zepdrix (zepdrix):

So let's first recognize the number of balls in the bag currently. 6 of one color, 10 of another. So there are a total of 16 balls in the bag right now. The ratio on the left, is the ratio of White balls to the entire number of balls. \[\large \frac{10}{16}=\frac{10+W}{16+112}\]On the right, the denominator represents the new number of total balls in the bag. W represents the number of WHITE balls that we should add to the bag to preserve this RATIO. Understand the setup? It's a little tricky :)

OpenStudy (anonymous):

a little can u sorta dumb the question down for me sorry

OpenStudy (anonymous):

?

zepdrix (zepdrix):

Maybe it'd be easier to think of them as percents. 10/16 is approximately 60%. So approximately 60% of the balls are white. If we add a crap load of balls to the bag, how many of them have to be white in order for the % of white balls to remain 60% of the new total number of balls. These are a little tough to setup. You have to write it two fractions. The ratio of the current balls, and set it equal to the number of balls you're looking for (our variable W) divided by the new number of balls.\[\large \frac{10}{16}=\frac{W}{128}\] With the way it's written here, W represents the number of White balls that should be in the bag now, so that the fractions 10/16 is equal to something/128. I'm not sure if there is an easier way to write it :( hmm.

OpenStudy (anonymous):



OpenStudy (anonymous):

heehee